What is the magnitude of the force on a charge of +2x10-7 C that is 0.3 m from a charge of -5x10-7 C?

1
Expert's answer
2022-03-28T14:24:50-0400

Force

"F=\frac{kq_1q_2}{r^2}"


"F=-\frac{9\times10^9\times2\times10^{-7}\times5\times10^{-7}}{0.3^2}=-0.01N"